OpenStudy (anonymous):

if 1/R=1/R_1+1/R_2+1/R_3, find dR/dR_1

OpenStudy (badhi):

do R_2,R_3 depend on R_1 if not it is like $$\frac{1}{y}=\frac{1}{x}+c$$ where y=R, x=R_1 and c=(1/R_2+1/R_3) (a constant). So just differentiate both side with respect to x (in this case R_1)